OpenStudy (anonymous):

determine the rational roots of 2x^5-11x^4+14x^3-2x^2+12x+9=0

OpenStudy (anonymous):

Hi, The first 3 roots are -.5, 3, and 3. . .____________________ -.5)2 . -11 . 14 . -2 . 12 . 9 . ____ -1__ 6_-10__ 6_-9_ . .2 . .-12 . 20 -12 . 18 . 0 .__________________ 3).2 . .-12 . 20 -12 . 18 ._____ 6_-18__ 6_-18_ . .2 . . -6 . .2 . .-6 . .0 . _______________ 3) .2 . -6 . .2 . .-6 . ____6__ 0___6_ . . 2 . .0 . . 2 . . 0 2x² + 2 = 0 2x² = -2 x² = -1 . . . . .__ x = ±√-1 x = ±i, -.5, 3, and 3 <==ANSWER I hope that helps!! :-)

OpenStudy (anonymous):

@adinarayan thank you!!!

OpenStudy (anonymous):

\[2 x^5-11 x^4+14 x^3-2 x^2+12 x+9=(x-3)^2 (2 x+1) \left(x^2+1\right) \]